OpenStudy (anonymous):

Medal + Fan + Testimony (5 questions)

OpenStudy (anonymous):

What is the solution to the system of equations represented by these two lines? A. (3, 1) B. (3, 0) C. (1, 3) D. (0, 6)

OpenStudy (anonymous):

it is (3,1) you are looking for the point where they intercept

OpenStudy (anonymous):

@ro561man Please don't give out the answer. :)

OpenStudy (anonymous):

@im.celibate The solution to the system of equations is where the lines intersect, or touch.

OpenStudy (anonymous):

COuld you help me with another one? @ro561man

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

What is the solution to the system of equations represented by the two equations?

OpenStudy (anonymous):

AHH! Sorry y = 2/3x y=2/3x + 4

OpenStudy (anonymous):

Im not sure about this one

OpenStudy (anonymous):

OK @AnswerMyQuestions

OpenStudy (anonymous):

We can graph them and see where the lines touch.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Here is the graph: https://www.desmos.com/calculator/2eaublufyg Where do the lines touch?

OpenStudy (anonymous):

they dont touch @AnswerMyQuestions

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Right, so do they have any solutions?

OpenStudy (igreen):

For there to be solutions the two lines must touch, and you say they don't, so..?

OpenStudy (anonymous):

WOuld it be 0?

OpenStudy (igreen):

Yep. No solutions.

OpenStudy (anonymous):

SO, 0,0? OKay..

OpenStudy (anonymous):

You were wrong -_-

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

Not (0,0)

OpenStudy (anonymous):

They don't have any solutions at all.

OpenStudy (anonymous):

So it isn't 0. It isn't anything.

OpenStudy (igreen):

No..I said NO SOLUTIONS. Not (0, 0).

OpenStudy (igreen):

(0, 0) IS a solution.

OpenStudy (igreen):

Understand? @im.celibate